OpenStudy (anonymous):

let f(x)=1-2x^2 and g(x) = x+1 compute the following (g o g)(-1)

OpenStudy (anonymous):

gg(-1)=1

OpenStudy (anonymous):

Steps: \[gg(x)=x+2\]\[gg(-1)=-1+2\]\[gg(-1)=1\]

OpenStudy (anonymous):

hay thanks it would be so easy if I understood it better thanks a ton

OpenStudy (anonymous):

I showed you about how a composite function comes, right? There're of the same concept.

OpenStudy (anonymous):

ok it is starting to get easier just remembering the order they go in is my problem.